OpenStudy (anonymous):

@ganeshie8 what would the other methods of manipulating exponents to rewrite them be? :)

ganeshie8 (ganeshie8):

First property : radical to exponent conversion (you already knw this) \(\large \sqrt[n]{a} = a^{\frac{1}{n}}\)

ganeshie8 (ganeshie8):

Second property : negative exponent property \(\large \dfrac{1}{x^{-n}} = x^n\)

ganeshie8 (ganeshie8):

Third property : multiplying two powers having the SAME BASE \(\large a^m \times a^n = a^{m+n} \) Fourth property : dividing two powers having the SAME BASE \(\large \dfrac{a^m}{ a^n} = a^{m\color{red}{-}n} \)

ganeshie8 (ganeshie8):

fine wid those ? give urself an example for each of them and convince urself :)

ganeshie8 (ganeshie8):

For ur previous problem, you need only the first two properties..

OpenStudy (anonymous):

okay yes:)  ? like this?